# Formal Degrees and Parabolic Induction: the Maximal Generic Case

### v2

@@ 3. Haar Measures
We fix a special maximal compact subgroup $K_{G}\subset G=\mathbf{G}(F)$, and for any closed subgroup $H$ of $G$, we write $K_{H}:=K_{G}\cap H$. In local harmonic analysis, one chooses the (left) Haar measure on any closed subgroup $H$ of $G$ such that $K_{H}$ has volume one. Thus if $H$ is a unimodular subgroup of $G$, $G/H$ is equipped with the left $G$-invariant quotient measure $d\dot{g}$, characterised by
@@
On the other hand, the formal degree conjecture uses another Haar meausre specified as follows. We have fixed a non-trivial additive character $\psi$ of $F$. Let $\mathcal{G}$ be a split form of $\mathbf{G}/\mathbf{A}_{\mathbf{G}}$ defined over $\mathfrak{o}$. Choose an $\overline{F}$-isomorphism
@@
and an invariant top form $\omega_{\mathcal{G}}$ on $\mathcal{G}$ defined over $\mathfrak{o}$ (which is equivalent to choosing a top form over $\mathfrak{o}$ on the Lie algebra of $\mathcal{G}$), with nonzero reduction (modulo $\varpi$). Clearly, $\omega_{\mathcal{G}}$ is unique up to a scalar in $\mathfrak{o}^{\times}$. Set $\omega_{\mathbf{G}}:=\eta^{*}(\omega_{\mathcal{G}})$ and let $dg:=dg_{\psi}$ be the Haar measure on $G/A_{\mathbf{G}}$ determined by $\omega_{\mathbf{G}}$ and the self-dual measure on $F$ with respect to $\psi$ (cf. §2.1). Then $dg$ depends only on $\psi$ and is independent of the choice of $\eta$ and $\omega_{\mathcal{G}}$ (cf. [Gro97][GG99]). We denote this Haar measure on $G/A_{\mathbf{G}}$ by $\mu_{2,\mathbf{G}}$.
@@
Let $\pi$ (resp. $\sigma$) be a discrete series representation of $G$ (resp. $M$). In order to use tools from harmonic analysis to deal with our problem, we have to first compute the quotient
@@
First we recall some basic facts about the motives associated to reductive groups introduced in [Gro97].
@@
$\mathfrak{M}_{\mathbf{H}}=\bigoplus_{d\geq 1}V_{d}(1-d),$
and one has
@@
Since $\mathbf{G}$ is unramified, there exists a connected reductive group scheme $\mathscr{G}$ over $\mathfrak{o}$ such that $\mathbf{G}$ is the generic fibre of $\mathscr{G}$ and $K_{G}:=\mathscr{G}(\mathfrak{o})$ is a hyperspecial maximal compact subgroup of $G=\mathbf{G}(F)$. In this case, it is also known that indeed $\mathscr{G}$ is also quasi-split over $\mathfrak{o}$ (cf. [Con14, §5.2.12]).
So we may fix a Borel subgroup $\mathbf{B}$ of $\mathbf{G}$ with its Levi decomposition $\mathbf{B}=\mathbf{T}\mathbf{U}$, such that $\mathbf{B},\mathbf{T},\mathbf{U}$ are defined over $\mathfrak{o}$. Let $\mathbf{T}_{0}$ be the maximal $F$-split subtorus of $\mathbf{T}$. Given $\alpha\in\Sigma(\mathbf{G},\mathbf{T}_{0})$, we denote by $\mathbf{U}_{\alpha}$ its (relative) root subgroup. This is a closed subgroup normalized by $\mathbf{T}_{0}$, and $\mathrm{Lie}(\mathbf{U}_{\alpha})$ is the direct sum of the root spaces of $\{\alpha,2\alpha\}\cap\Phi(\mathbf{G},\mathbf{T}_{0})$. Then the subgroup $I_{G}$ of $G$ generated by
@@
where $a(\mathfrak{M}_{\mathbf{G}})$ is the Artin conductor of $\mathfrak{M}_{\mathbf{G}}$, $I_{F}$ is the inertia group of $F$ and $F^{\mathrm{ur}}$ the maximal unramified extension of $F$ in $\overline{F}$ (cf. [GG99, §4, §5]). Since $\mathbf{G}$ is unramified,
@@
$\nu_{\mathbf{G}}(I_{G}^{+})=q^{-\frac{1}{2}(\dim\mathbf{G}-\dim\mathbf{T})}.$
@@
$\nu_{\mathbf{M}}(I_{M}^{+})=q^{-\frac{1}{2}(\dim\mathbf{M}-\dim\mathbf{T})},$
@@
Next for any $\alpha\in\Sigma(\mathbf{G},\mathbf{T}_{0})$, take a preimage $\tilde{\alpha}\in\Sigma(\mathbf{G},\mathbf{T})$ of $\alpha$ in the absolute root system. Recall that the Galois group $\Gamma=\mathrm{Gal}(\overline{F}/F)$ acts on the absolute root datum. Let $\Gamma_{\tilde{a}}\subset\Gamma$ be the stabilizer of $\tilde{\alpha}$ and $F_{\tilde{\alpha}}\subset\overline{F}$ be the fixed field. Then the natural $F_{\tilde{\alpha}}$-embedding $\mathbf{U}_{\tilde{\alpha}}\hookrightarrow\mathbf{U}_{\alpha}$ induces an $F$-isomorphism
(cf. [KP23, pp. 97]). In our setting, $\mathbf{U}_{\alpha}$ and the above isomorphism are defined over $\mathfrak{o}$. We denote by $\mathfrak{o}_{\tilde{\alpha}}$ the ring of integers of $F_{\tilde{\alpha}}$, with the maximal ideal $\mathfrak{p}_{\tilde{\alpha}}$. Also, $F_{\tilde{\alpha}}/F$ is unramified, so that
$[U_{\alpha}\cap K_{G}:U_{\alpha}\cap I_{G}^{+}]$ $=[\mathbf{U}_{\alpha}(\mathfrak{o}):\mathbf{U}_{\alpha}(\mathfrak{p})]$
$=[\mathbf{U}_{\tilde{\alpha}}(\mathfrak{o}_{\tilde{\alpha}}):\mathbf{U}_{\tilde{\alpha}}(\mathfrak{p}_{\tilde{\alpha}})]$
$=q^{[F_{\tilde{\alpha}}:F]}=q^{\dim\mathbf{U}_{\alpha}}\quad(\alpha\in\Sigma_{\mathrm{red}}^{-}(\mathbf{G},\mathbf{T}_{0})).$
@@
(cf. [Wal03, pp. 241]). Since $\mathrm{Lie}(\mathbf{N})=\bigoplus_{\alpha\in\Sigma_{\mathrm{red}}(\mathbf{P})}\mathrm{Lie}(\mathbf{U}_{\alpha})$, $\sum_{\alpha\in\Sigma_{\mathrm{red}}(\mathbf{P})}\dim\mathbf{U}_{\alpha}=\dim\mathbf{N}$,
@@
then the above proposition doesn’t hold in general. (For some examples, cf. Appendix E of [GI14], the tables of case A.) For this reason, we add the unramified condition for our main theorem to make the explicit formula less technical.
@@ 4.1. Results from Langlands-Shahidi method
be the corresponding fundamental weight. Denote by $r$ the adjoint representation of ${{}^{L}\!M}$ on ${{}^{L}\mathfrak{n}}:=\mathrm{Lie}({{}^{L}\!N})$. Then $r$ decomposes as $(r,{{}^{L}\mathfrak{n}})=\bigoplus_{i=1}^{m}(r_{i},V_{i})$, where each irreducible component $(r_{i},V_{i})$ is given by
@@
cf. [Sha88, Proposition 4.1]. Let $w_{0}\in W=W^{\mathbf{G}}(\mathbf{T}_{0})$ be the element of the Weyl group characterized by $w_{0}(\Delta-\{\alpha\})\subset\Delta$ and $w_{0}\alpha\in\Sigma^{-}$, i.e. the longest element modulo $W^{\mathbf{M}}(\mathbf{T}_{0})$.
@@
Theorem 4.1 ([Sha90, Theorem 8.1]).
@@
Since by [Sha90, Proposition 7.8] we have (for any irreducible generic supercuspidal $\sigma$)
@@
for $\sigma^{\prime}$ in a Zariski open subset of $\mathcal{O}$. Note that the definition here is compatible with [Hei04, §1.5] but differs from that in [Sha90][Wal03]. When $\mathbf{P}=\mathbf{M}\mathbf{N}$ is maximal, the $\mu$-function here is equal to $\gamma(\mathbf{G}/\mathbf{M})\mu$ defined in [Sha90][Wal03].
The main theorem together with (3.12) of [Sha90] can then be formulated as
@@
where $\gamma^{\mathrm{Sh}}$ are Shahidi’s $\gamma$-factors characterized in [Sha90, Theorem 3.5].

### v3

@@ 3. Haar Measures
Fix a Haar measure $\mu$ on $G/A$. Let $(\pi,V_{\pi})$ be a discrete series of $G$ and (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) a $G$-invariant inner product on $V_{\pi}$. Recall that by the Schur orthogonality relation of matrix coefficients, there exists a unique positive real number $d(\pi)=d(\pi,\mu)\in\mathbb{R}^{\times}_{>0}$, called the formal degree of $\pi$ with respect to $\mu$, such that
Clearly, $d(\pi)=d(\pi,\mu)$ is independent of the choice of (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) but depends on the choice of $\mu$, and $d(\pi,\lambda\mu)=\lambda^{-1}\cdot d(\pi,\mu)$ for $\lambda\in\mathbb{R}^{\times}_{>0}$.
We fix a special maximal compact subgroup $K_{G}\subset G=\mathbf{G}(F)$, and for any closed subgroup $H$ of $G$, we write $K_{H}:=K_{G}\cap H$. In local harmonic analysis, one chooses the (left) Haar measure on any closed subgroup $H$ of $G$ such that $K_{H}$ has volume one. Thus if $H$ is a unimodular subgroup of $G$, $G/H$ is equipped with the left $G$-invariant quotient measure $d\dot{g}$, characterized by
@@
On the other hand, the formal degree conjecture uses another Haar measure specified as follows. We have fixed a non-trivial additive character $\psi$ of $F$. Let $\mathcal{G}$ be a split form of $\mathbf{G}/\mathbf{A}_{\mathbf{G}}$ defined over $\mathfrak{o}=\mathfrak{o}_{F}$. Choose an $\overline{F}$-isomorphism
@@
and an invariant top form $\omega_{\mathcal{G}}$ on $\mathcal{G}$ defined over $\mathfrak{o}$ (which is equivalent to choosing a top form over $\mathfrak{o}$ on the Lie algebra of $\mathcal{G}$), with nonzero reduction (modulo $\varpi=\varpi_{F}$). Clearly, $\omega_{\mathcal{G}}$ is unique up to a scalar in $\mathfrak{o}^{\times}$. Set $\omega_{\mathbf{G}}:=\eta^{*}(\omega_{\mathcal{G}})$ and let $dg:=dg_{\psi}$ be the Haar measure on $G/A_{\mathbf{G}}$ determined by $\omega_{\mathbf{G}}$ and the self-dual measure on $F$ with respect to $\psi$ (cf. §2.1). Then $dg$ depends only on $\psi$ and is independent of the choice of $\eta$ and $\omega_{\mathcal{G}}$ (cf. [gross-motive][gan-gross-haar-measure]). We denote this Haar measure on $G/A_{\mathbf{G}}$ by $\mu_{2,\mathbf{G}}$.
@@
Let $\pi$ (resp. $\sigma$) be a discrete series representation of $G$ (resp. $M$). In order to use tools from harmonic analysis to deal with our problem, we have to first evaluate the quotient
@@
First we recall some basic facts about the motives associated to reductive groups introduced in [gross-motive].
@@
$\mathfrak{M}_{\mathbf{H}}=\bigoplus_{d\geq 1}V_{d}(1-d)$
where $V_{d}(k)$ is the $k$-Tate twist of $V_{d}$, and one has
@@
Since $\mathbf{G}$ is unramified, there exists a connected reductive group scheme $\mathscr{G}$ over $\mathfrak{o}$ such that $\mathbf{G}$ is the generic fibre of $\mathscr{G}$ and $K_{G}:=\mathscr{G}(\mathfrak{o})$ is a hyperspecial maximal compact subgroup of $G=\mathbf{G}(F)$. In this case, it is also known that indeed $\mathscr{G}$ is also quasi-split over $\mathfrak{o}$ (cf. [conrad-reductive-group-schemes, §5.2.12]).
So we may fix a Borel subgroup $\mathbf{B}$ of $\mathbf{G}$ with its Levi decomposition $\mathbf{B}=\mathbf{T}\mathbf{U}$, such that $\mathbf{B},\mathbf{T},\mathbf{U}$ are defined over $\mathfrak{o}$. Let $\mathbf{T}_{0}$ be the maximal $F$-split subtorus of $\mathbf{T}$. Given $\alpha\in\Sigma_{\mathrm{red}}(\mathbf{G},\mathbf{T}_{0})$, we denote by $\mathbf{U}_{\alpha}$ its (relative) root subgroup. This is a closed subgroup normalized by $\mathbf{T}_{0}$, and $\mathrm{Lie}(\mathbf{U}_{\alpha})$ is the direct sum of the root spaces of $\{\alpha,2\alpha\}\cap\Sigma(\mathbf{G},\mathbf{T}_{0})$. Then the subgroup $I_{G}$ of $G$ generated by
@@
where $a(\mathfrak{M}_{\mathbf{G}})$ is the Artin conductor of $\mathfrak{M}_{\mathbf{G}}$, $I_{F}$ is the inertia group of $F$ and $F^{\mathrm{ur}}$ the maximal unramified extension of $F$ in $\overline{F}$ (cf. [gross-motive, pp. 295] and [gan-gross-haar-measure, §5]). Since $\mathbf{G}$ is unramified,
@@
$\nu_{\mathbf{G}}(I_{G}^{+})=q^{-\frac{1}{2}(\dim\mathbf{G}+\dim\mathbf{T})}.$
@@
$\nu_{\mathbf{M}}(I_{M}^{+})=q^{-\frac{1}{2}(\dim\mathbf{M}+\dim\mathbf{T})},$
@@
Next for any $\alpha\in\Sigma_{\mathrm{red}}(\mathbf{G},\mathbf{T}_{0})$, we have
Fact 3.2.
$[\mathbf{U}_{\alpha}(\mathfrak{o}):\mathbf{U}_{\alpha}(\mathfrak{p})]=q^{\dim\mathbf{U}_{\alpha}}$.
Let us just remark here that this follows directly from the structure of the relative root subgroups $\mathbf{U}_{\alpha}$ and the unramified condition. We shall write down the related details as a proof later.
With this fact, for $\alpha\in\Sigma_{\mathrm{red}}^{-}(\mathbf{G},\mathbf{T}_{0})$ we have
$[U_{\alpha}\cap K_{G}:U_{\alpha}\cap I_{G}^{+}]=[\mathbf{U}_{\alpha}(\mathfrak{o}):\mathbf{U}_{\alpha}(\mathfrak{p})]=q^{\dim\mathbf{U}_{\alpha}}.$
@@
(cf. [waldspurger-plancherel, pp. 241]). Since $\mathrm{Lie}(\overline{\mathbf{N}})=\bigoplus_{\alpha\in\Sigma_{\mathrm{red}}(\overline{\mathbf{P}})}\mathrm{Lie}(\mathbf{U}_{\alpha})$ and $\sum_{\alpha\in\Sigma_{\mathrm{red}}(\overline{\mathbf{P}})}\dim\mathbf{U}_{\alpha}=\dim\overline{\mathbf{N}}=\dim\mathbf{N}$, we get
@@
Now we recollect the related details of the above fact. The reader is referred to [kaletha, pp. 96-97] for an excellent and clear explanation of the structure of relative root subgroups of quasi-split groups. With this the fact is nothing more than a direct consequence of the unramified condition.
Proof of Fact 3.2.
Let $\alpha\in\Sigma_{\mathrm{red}}(\mathbf{G},\mathbf{T}_{0})$. There are two cases. We take a preimage $\tilde{\alpha}\in\Sigma(\mathbf{G},\mathbf{T})$ of $\alpha$ in the absolute root system. Recall that the Galois group $\Gamma=\mathrm{Gal}(\overline{F}/F)$ acts on $\Sigma(\mathbf{G},\mathbf{T})$. Let $\Gamma_{\tilde{\alpha}}\subset\Gamma$ be the stabilizer of $\tilde{\alpha}$ and $F_{\tilde{\alpha}}\subset\overline{F}$ be the fixed field, so that $F_{\tilde{\alpha}}/F$ is unramified by our assumption. We then denote by $\mathfrak{o}_{\tilde{\alpha}}$ the ring of integers of $F_{\tilde{\alpha}}$, with the maximal ideal $\mathfrak{p}_{\tilde{\alpha}}$.
  • First, suppose $2\alpha\notin\Sigma(\mathbf{G},\mathbf{T}_{0})$. In this case the natural $F_{\tilde{\alpha}}$-embedding $\mathbf{U}_{\tilde{\alpha}}\hookrightarrow\mathbf{U}_{\alpha}$ induces an $F$-isomorphism
    (cf. [kaletha, pp. 97]). In our setting, $\mathbf{U}_{\alpha}$ and the above isomorphism are defined over $\mathfrak{o}$. Since $F_{\tilde{\alpha}}/F$ is unramified,
    $[\mathbf{U}_{\alpha}(\mathfrak{o}):\mathbf{U}_{\alpha}(\mathfrak{p})]$ $=[\mathbf{U}_{\tilde{\alpha}}(\mathfrak{o}_{\tilde{\alpha}}):\mathbf{U}_{\tilde{\alpha}}(\mathfrak{p}_{\tilde{\alpha}})]$
    $=[\mathfrak{o}_{\tilde{\alpha}}:\mathfrak{p}_{\tilde{\alpha}}]=q^{[F_{\tilde{\alpha}}:F]}=q^{\dim\mathbf{U}_{\alpha}}.$
  • The second situation is $2\alpha\in\Sigma(\mathbf{G},\mathbf{T}_{0})$. In this case the structure of the relative root subgroup $\mathbf{U}_{\alpha}$ is more complicated to explain. Indeed this case only occurs when $\Sigma(\mathbf{G},\mathbf{T})$ has an irreducible factor of type $A_{2n}$ and $\Gamma$ acts nontrivially on this type $A_{2n}$ factor.
    The preimage of $\alpha$ in $\Sigma(\mathbf{G},\mathbf{T})$ is of even order, and there is a unique $\tilde{\alpha}^{\prime}$ in the preimage which is not orthogonal to $\tilde{\alpha}$. Then $\tilde{\beta}:=\tilde{\alpha}+\tilde{\alpha}^{\prime}$ lies in $\Sigma(\mathbf{G},\mathbf{T})$ and is mapped to $2\alpha\in\Sigma(\mathbf{G},\mathbf{T}_{0})$. We denote by $F_{\tilde{\beta}},\mathfrak{o}_{\tilde{\beta}},\mathfrak{p}_{\tilde{\beta}}$ the similar notations for $\tilde{\beta}$ (as those for $\tilde{\alpha}$); then $F_{\tilde{\alpha}}/F_{\tilde{\beta}}$ is quadratic. We have $\mathbf{U}_{\alpha}\simeq\mathrm{Res}_{F_{\tilde{\beta}}/F}\mathbf{U}_{[\tilde{\alpha}]}$ over $\mathfrak{o}$, where $\mathbf{U}_{[\tilde{\alpha}]}$ is isomorphic to the unipotent radical of a Borel subgroup of the quasi-split $\mathrm{SU}_{3}(F_{\tilde{\alpha}}/F_{\tilde{\beta}})$ ([kaletha, pp. 97]). More explicitly, if we denote by $a\mapsto\overline{a}$ the nontrivial element of $\mathrm{Gal}(F_{\tilde{\alpha}}/F_{\tilde{\beta}})$, then
    $\mathbf{U}_{[\tilde{\alpha}]}=\{u(a,b):=\begin{bmatrix}1&a&b\\ &1&\overline{a}\\ &&1\end{bmatrix}\mid a,b\in F_{\tilde{\alpha}},~{}a\overline{a}=b+\overline{b}\}$
    as a group over $F_{\tilde{\beta}}$. We have an exact sequence
    $1\rightarrow\mathbf{U}^{0}_{[\tilde{\alpha}]}\rightarrow\mathbf{U}_{[\tilde{\alpha}]}\rightarrow\mathrm{Res}_{F_{\tilde{\alpha}}/F_{\tilde{\beta}}}\mathbf{G}_{a}\rightarrow 1$
    of $F_{\tilde{\beta}}$-groups, where $\mathbf{G}_{a}$ is the additive group of $F_{\tilde{\alpha}}$ and
    $\mathbf{U}^{0}_{[\tilde{\alpha}]}=\{u(0,b)\mid b\in F_{\tilde{\alpha}},~{}b+\overline{b}=0\}.$
    Since $F_{\tilde{\alpha}}\supset F_{\tilde{\beta}}\supset F$ are unramified, $\mathbf{U}^{0}_{[\tilde{\alpha}]}(\mathfrak{o}_{\tilde{\beta}})/\mathbf{U}^{0}_{[\tilde{\alpha}]}(\mathfrak{p}_{\tilde{\beta}})$ is just the subgroup of $\mathfrak{o}_{\alpha}/\mathfrak{p}_{\alpha}$ consisting of elements of quadratic trace 00, hence of order $[\mathfrak{o}_{\tilde{\beta}}:\mathfrak{p}_{\tilde{\beta}}]$, and
    $[\mathbf{U}_{\alpha}(\mathfrak{o}):\mathbf{U}_{\alpha}(\mathfrak{p})]$ $=[\mathbf{U}_{[\tilde{\alpha}]}(\mathfrak{o}_{\tilde{\beta}}):\mathbf{U}_{[\tilde{\alpha}]}(\mathfrak{p}_{\tilde{\beta}})]=[\mathfrak{o}_{\tilde{\beta}}:\mathfrak{p}_{\tilde{\beta}}]^{3}$
    $=q^{[F_{\tilde{\beta}}:F]\cdot\dim_{F_{\tilde{\beta}}}(U_{[\tilde{\alpha}]})}=q^{\dim\mathbf{U}_{\alpha}}.$
@@
then the above proposition does not hold in general. (For some examples, cf. Appendix E of [gan-ichino-formal-degree], the tables of case A.) For this reason, we add the unramified condition for our main theorem to make the explicit formula less technical.
@@ 4.1. Results from Langlands-Shahidi method
be the corresponding fundamental weight. Denote by $r$ the adjoint representation of ${{}^{L}\!M}$ on $\widehat{\mathfrak{n}}:=\mathrm{Lie}(\widehat{N})$. Then $r$ decomposes as $(r,\widehat{\mathfrak{n}})=\bigoplus_{i=1}^{m}(r_{i},V_{i})$, where each irreducible component $(r_{i},V_{i})$ is given by
@@
cf. [shahidi-ramanujan, Proposition 4.1]. Let $W=W^{\mathbf{G}}(\mathbf{T}_{0})$ be the Weyl group of $\mathbf{G}$ with respect to the maximal split torus $\mathbf{T}_{0}$ and $w_{0}\in W=W^{\mathbf{G}}(\mathbf{T}_{0})$ be the element of the Weyl group characterized by $w_{0}(\Delta-\{\alpha\})\subset\Delta$ and $w_{0}\alpha\in\Sigma^{-}$, i.e. the longest element modulo the Weyl group of $\mathbf{M}$.
@@
Theorem 4.1 ([shahidi-plancherel, Theorem 8.1]).
@@
Since by [shahidi-plancherel, Proposition 7.8] we have (for any irreducible generic supercuspidal $\sigma$)
@@
for $\sigma^{\prime}$ in a Zariski open subset of $\mathcal{O}$. Note that the definition here is compatible with [heiermann-spectrales, §1.5] but differs from that in [shahidi-plancherel][waldspurger-plancherel]. When $\mathbf{P}=\mathbf{M}\mathbf{N}$ is maximal, the $\mu$-function here is equal to $\gamma(\mathbf{G}/\mathbf{M})^{-2}\mu$ defined in [shahidi-plancherel][waldspurger-plancherel].
The main theorem 3.5 and corollary 3.6 of [shahidi-plancherel] then reads
@@
where $\gamma^{\mathrm{Sh}}$ are Shahidi’s $\gamma$-factors characterized in [shahidi-plancherel, Theorem 3.5].